OpenStudy (anonymous):

write in set builder notation even numbers between 27 and 39 PLEASE HELP!

OpenStudy (debbieg):

I think you're looking for something like....\[ \left\{ x|x \in \mathbb{R}, 27<x<39,x=2k, k \in \mathbb{N} \right\}\]...maybe??

OpenStudy (debbieg):

Could use \[k \in \mathbb{Z}\] also, same resulting set. :)

OpenStudy (anonymous):

THANK YOU ABSOLUTELY SOOOO MUCH i've been waiting for someone to answer my question for an hour!

OpenStudy (debbieg):

Happy to help. :)